OpenStudy (foreverandalways13):

im not understanding slope intercept equations....anyone that can help???

OpenStudy (anonymous):

where are u confused?

OpenStudy (galleara):

y=mx+b m is the slope you are given (we'll say 4) and if you are given a point say (1,2) y=2 and x=1 entered into the equation you would have this: 2=4(1)+b where you are looking for b. after multiplying 4*1, subtract it from 2 and you will get b=-2. Then your equation will look like y=4x-2. Does this help?

OpenStudy (anonymous):

cool..

OpenStudy (foreverandalways13):

ok lets say they want us to solve for x....how would you do that?

OpenStudy (foreverandalways13):

but yes it helps alot!!! thanks!

OpenStudy (galleara):

well if you are solving for x, using the same numbers as above, we would put in y=mx+b 2=4x-2 add 2+2 --> 4=4x divde by 4 and you get 1

OpenStudy (foreverandalways13):

oh that makes sense!!! thanks again!

OpenStudy (galleara):

you're welcome :)